Consider two Tanzanian Banks, A and B, with the following quotes which are expressed as Ksh/Tsh.

Bank A Bank B
Tsh quote Bid Ask Bid Ask
Ksh0.06 Ksh.0.07 Ksh 0.08 Ksh0.09
Required
Compute the gain from locational arbitrage to a Tanzanian investor with Kshs.30 million to invest (5 marks)
1
Expert's answer
2019-11-29T10:03:20-0500

Gain from locational arbitrage is realized by buying form a bank whose ask price (A-0.07) is lower than the bid price of the other (B-0.08).Bank A buys 1 ksh at

"\frac{1}{0.07}=Tsh. 14.29"

.

Ksh. 30 million is bought at



"30000000\times14.29=Tsh.428700000"

Bank B sells 1 Tsh. at ksh. 0.08

Tsh. 428700000 are sold for


"428700000\times 0.08=ksh.34296000"

Gain from trade is therefore,


"34296000-30000000=ksh. 4296000"